What principles can you use to find the spe

What principles can you use to find the speed of the block at the base of the incline? (choose all that apply)

___ conservation of linear momentum

___ conservation of mechanical energy

___ kinematics
Answer #1

The speed of the block at the base of the incline can be found by using the principle of conservation of mechanical energy.

At the top of the incline total mechanical energy of the block is its potential energy. (initial speed of block at the top is zero, so kinetic energy is zero). that is total mechanical energy at the top of incline = mgh

At the bottom of the incline total mechanical energy of the block is its kinetic energy.(height=0 , potential energy=0)

that is total mechanical energy at the bottom of incline is =\frac{1}{2}mv^2

conserving mechanical energy of the block

mgh=\frac{1}{2}mv^2

v=\sqrt{2gh}